Why the correct options are B and D only, not including C and E?
QUESTION 1 Anita has $600 per month to spend on electricity and other goods. Price of other goods is one, while price of electricity is 50.2 per kilowatt hour (kwH) for the first 1,000 kwH of power per month and 50.04 for any additional kwH. Select all that applies: O a. For the amount of electricity less than 1,000 kwH per month, the slope of her budget line is-0.1. b. For the amount of electricity less than 1,000 kWH per month, the slope of her budget line is-02. OC. For the amount of electricity more than 1,000 kwH per month, the slope of her budget line is-0.24 O d. A bundle of 5,000 kwH and 240 of other goods is feasible for her. O e. A bundle of 2,000 kwH and 300 of other goods is feasible for her. t. A bundle of 6,000 kwH and 250 of other goods is feasible for her.
Answer #1

Answer : The answers are b and d.

The slope of budget line for less than 1,000 kwH = - (price of electricity / price of other goods) = - (0.2 / 1)

Therefore, The slope of budget line = - 0.2

Thus, option b is correct.

For first 1,000 kwH the price is $0.2. So for first 1,000 kwH the spending amount = 1000 * 0.2 = $200

For other 4,000 kwH the price is $0.04. So, for this additional 4,000 kwH the spending amount = 4000 * 0.04 = $160

Total spending amount on 5,000 kwH = 200 + 160 = $360.

Remaining spending amount for Anita = 600 - 360 = $240. This $240 remaining spending amount will be spend on other goods.

Thus, option d is correct.

In case of option c, the price of electricity per kwH is 0.04. So, in this case the slope of budget line for more than 1,000 kwH = - (price of electricity / price of other goods) = - 0.04 / 1 = - 0.04 . Hence option c is not included.

In case of option e, for first 1,000 kwH the price is $0.2. So for first 1,000 kwH the spending amount = 1000 * 0.2 = $200. For additional 1,000 kwH the price is $0.04. So for this additional 1,000 kwH the spending amount = 1000 * 0.04 = $40

Total spending amount on 2,000 kwH = 200 + 40 = $240

Remaining spending amount for Anita = 600 - 240 = $360. This $360 remaining spending amount will be spend on other goods. Hence option e is not included.